AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 08:28:18 2025
Author
*
:
Subject
*
:
<p>11:01</p> <p>Run 35 started cooled beam on target</p> <p>Detectors IN (left arm pot =36.5 mm, right arm pot = 70.49 mm)</p> <p>Electron cooler voltage 751.4 this corresponds to 1.3436 MeV/u. This should be on the peak of the resonance.</p> <p>Target density is 3.75e11 at/cm2, ring pressure looks OK YR09 1.43e-10 mbar.</p> <p>Horizontal beam position is -7.5 mm, vertical beam position is 0.0 mm. Same as run 33.</p> <p>Beam is in the ring for 8 seconds and total cycle time between injections is 12 seconds. Same as run 33</p> <p>Beam injection looks stable. ~2e7 injected and no missed injections so far.</p> <p>Merger OK, Tape server OK</p> <p>Merger statistics zero'd</p> <p>All histograms and statistics zero'd, monitoring code reset.</p> <p>ADC data item statistics looks as before, very few counts until injection into the ring. OK</p> <p>FEE temperatures look OK</p> <p>System checks OK</p> <p>Leakage current trending upwards, same as previous elog.</p> <p>13:33</p> <p>FEE temperatures OK - all FEEs at least 1.5C below vertex warning</p> <p>ADC data items statistics OK</p> <p>Merger and Tape server OK</p> <p>Target density 3.8e11 at/cm2, ring pressure stable.</p> <p>Ring injections look very stable</p> <p>Leakage current now decreasing (attachment 1)</p> <p>Monitoring code xy and energy histograms (attachment 2+3)</p> <p>15:03</p> <p>System wide checks ok. Merger ok. Temperatures ok.</p> <p>Attatchment 4: Stats</p> <p>Attatchment 5: Temperatures</p> <p>Attatchments 6-9: Waveforms.</p> <p>Attatchment 10: Leakage current</p> <p>Attatchment 11: lowenergyEytotal plots</p> <p>17:09</p> <p>System wide check ok. Merger ok. Temperature ok.</p> <p>Attachment 12: Stats</p> <p>Attachment 13: Temperatures</p> <p>Attachment 14-18: Waveforms (Layout 7: strange behaviour in aida01, went to ASIC Control and did "Check ASIC control" three times, which resolved the problem)</p> <p>Attachment 19: Leakage current</p> <p>Attachment 20: lowenergyEytotal plots</p> <p>Gas pressure was going up a little bit: changed Y position of the nozzle from 1000 to 950.</p> <p>17:55</p> <p>The gas jet target pressure started to fall sharply at 17:34 and went down to zero within 2 minutes.</p> <p>It came back while it we were changing the nozzle position but at 17:49 fell to zero again.</p> <p>Currently waiting for Nikos to came and have look at it.</p> <p>18:10</p> <p>Target back in place and has a density of 4.04e11. Beam transformer looks similar to pre plummeting target density.</p> <p> </p> <p> </p> <p>18.55 Check ASIC load FEE64 aida16 asic#3 HEC channels active</p> <p> Following check load merger rate decreased significantly.</p> <p>19:14</p> <p>Gas density remains stable around 4e11 after the incident at 17:34 and the beam transformer singal looks ok too. Gas pressure ok.</p> <p>Merger: All aidas turn green but the merger has decreased significantly (too about 306 000 items/sec) as mentioned above.</p> <p>System wide check ok. Temperature ok.</p> <p>Attachment 21: Stats</p> <p>Attachment 22: Temperatures</p> <p>Attachment 23-26: Waveforms</p> <p>Attachment 27: Leakage current</p> <p> </p> <p> </p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown